# Is ChatGPT the Future of Causal Text Mining? A Comprehensive Evaluation and Analysis

### 22 Feb 2024 v1

@@ Abstract
Causality is fundamental in human cognition and has drawn attention in diverse research fields. With growing volumes of textual data, discerning causalities within text data is crucial, and causal text mining plays a pivotal role in extracting meaningful patterns. This study conducts comprehensive evaluations of ChatGPT’s causal text mining capabilities. Firstly, we introduce a benchmark that extends beyond general English datasets, including domain-specific and non-English datasets. We also provide an evaluation framework to ensure fair comparisons between ChatGPT and previous approaches. Finally, our analysis outlines the limitations and future challenges in employing ChatGPT for causal text mining. Specifically, our analysis reveals that ChatGPT serves as a good starting point for various datasets. However, when equipped with a sufficient amount of training data, previous models still surpass ChatGPT’s performance. Additionally, ChatGPT suffers from the tendency to falsely recognize non-causal sequences as causal sequences. These issues become even more pronounced with advanced versions of the model, such as GPT-4. In addition, we highlight the constraints of ChatGPT in handling complex causality types, including both intra/inter-sentential and implicit causality. The model also faces challenges with effectively leveraging in-context learning and domain adaptation. Our code is available on https://github.com/retarfi/gemcausal
@@ 3.1 Datasets
For comprehensive evaluations of CTM, we organize a variety of datasets related to causal text mining across different domains, and languages.
@@ Appendix A Datasets Details
The following sections detail the causal text mining dataset utilized in our study: AltLex (Hidey and McKeown, 2016) investigates causal relations in English Wikipedia articles, focusing on alternative lexicalization connectives (AltLex) within a single sentence. These AltLex connectives appear in a broader range of linguistic forms compared to explicit markers, exemplified by phrases such as “This may help explain why” and “This activity produced.” However, one limitation of this dataset is its assumption that all words preceding and succeeding the causal cues are the sole spans corresponding to the cause and effect intended for extraction.
BECauSE V2.0 (BECauSE) (Dunietz et al., 2017) is designed to annotate intra-sentential explicit causal relations, aiming to encapsulate the diverse constructions utilized to convey cause and effect. The arguments directed to causal relation instances are phrase units. The dataset is derived from several sources, namely the New York Times Annotated Corpus (NYT) (NYT), Penn Treebank (PTB) (Marcus-94), Congressional Hearings of the 2014 NLP Unshared Task in PoliInformatics (CHRG) (Smith-14), and the Manually Annotated Sub-Corpus (MASC) (MASC).
Causal-TimeBank (CTB) (Mirza and Tonelli, 2014; Mirza et al., 2014) and EventStoryLine (ESL) (Caselli and Vossen, 2016, 2017) datasets are widely recognized in the field of Event Causality Identification (ECI), which concentrates on discerning causal links between events within textual data. For instance, the ECI model identifies a causal link between “earthquake” and “tsunami” in the sentence “The earthquake generated a tsunami.” The CTB dataset originates from the TimeBank corpus of the TempEval-3 task (uzzaman-etal-2013-tempeval3) and is designed to annotate solely explicit causal relations via a rule-driven algorithm. In contrast, ESL derives from an extended version of the EventCorefBank (ECB+) (cybulska-vossen-2014-using) and encompasses explicit and implicit causality. Both datasets address intra-sentential as well as inter-sentential causality. However, a limitation is evident in that only the initial words of an event are tagged, leading to the omission of the context from the extracted arguments.

### 23 Feb 2024 v2

@@ Abstract
Causality is fundamental in human cognition and has drawn attention in diverse research fields. With growing volumes of textual data, discerning causalities within text data is crucial, and causal text mining plays a pivotal role in extracting meaningful patterns. This study conducts comprehensive evaluations of ChatGPT’s causal text mining capabilities. Firstly, we introduce a benchmark that extends beyond general English datasets, including domain-specific and non-English datasets. We also provide an evaluation framework to ensure fair comparisons between ChatGPT and previous approaches. Finally, our analysis outlines the limitations and future challenges in employing ChatGPT for causal text mining. Specifically, our analysis reveals that ChatGPT serves as a good starting point for various datasets. However, when equipped with a sufficient amount of training data, previous models still surpass ChatGPT’s performance. Additionally, ChatGPT suffers from the tendency to falsely recognize non-causal sequences as causal sequences. These issues become even more pronounced with advanced versions of the model, such as GPT-4. In addition, we highlight the constraints of ChatGPT in handling complex causality types, including both intra/inter-sentential and implicit causality. The model also faces challenges with effectively leveraging in-context learning and domain adaptation. We release our code to support further research and development in this field.
@@ 3.1 Datasets
For comprehensive evaluations of CTM, we organize a variety of datasets related to causal text mining across different domains and languages.
@@ Appendix A Datasets Details
This section details the causal text mining dataset utilized in our study. AltLex (Hidey and McKeown, 2016) investigates causal relations in English Wikipedia articles, focusing on alternative lexicalization connectives (AltLex) within a single sentence. These AltLex connectives appear in a broader range of linguistic forms compared to explicit markers, exemplified by phrases such as “This may help explain why” and “This activity produced.” However, one limitation of this dataset is its assumption that all words preceding and succeeding the causal cues are the sole spans corresponding to the cause and effect intended for extraction.
BECauSE V2.0 (BECauSE) (Dunietz et al., 2017) is designed to annotate intra-sentential explicit causal relations, aiming to encapsulate the diverse constructions utilized to convey cause and effect. The arguments directed to causal relation instances are phrase units. The dataset is derived from several sources, namely the New York Times Annotated Corpus (NYT) (Sandhaus, 2008), Penn Treebank (PTB) (Marcus et al., 1994), Congressional Hearings of the 2014 NLP Unshared Task in PoliInformatics (CHRG) (Smith et al., 2014), and the Manually Annotated Sub-Corpus (MASC) (Ide et al., 2010).
Causal-TimeBank (CTB) (Mirza and Tonelli, 2014; Mirza et al., 2014) and EventStoryLine (ESL) (Caselli and Vossen, 2016, 2017) datasets are widely recognized in the field of Event Causality Identification (ECI), which concentrates on discerning causal links between events within textual data. For instance, the ECI model identifies a causal link between “earthquake” and “tsunami” in the sentence “The earthquake generated a tsunami.” The CTB dataset originates from the TimeBank corpus of the TempEval-3 task (UzZaman et al., 2013) and is designed to annotate solely explicit causal relations via a rule-driven algorithm. In contrast, ESL derives from an extended version of the EventCorefBank (ECB+) (Cybulska and Vossen, 2014) and encompasses explicit and implicit causality. Both datasets address intra-sentential as well as inter-sentential causality. However, a limitation is evident in that only the initial words of an event are tagged, leading to the omission of the context from the extracted arguments.